Grade 11Algebra

If the tangent to the circle x2 + y2 = r2 at the point (a, b) meets the coordinate axes at the point A, B and O is the origin then the area of the triangle OAB is ( 1) r5/ ab (2) r2 / ab (3) r4 / 2ab (4) r2 / 2ab

Nitin Kumar
9 Years agoGrade 11

1 Answer

Atharv
6 Years ago
Radius=√a2+b2
Equantion of AB =ax+by=r2
Or equation if AB=x/r2/a +y/r2/b=1
OA=r2/a.   OB=r2/b
Area=1/2*OA*OB =r4/2ab